# UI-AGILE: Advancing GUI Agents with Effective Reinforcement Learning and Precise Inference-Time Grounding

### v2

@@ Abstract
The emergence of Multimodal Large Language Models (MLLMs) has driven significant advances in Graphical User Interface (GUI) agent capabilities. Nevertheless, existing GUI agent training and inference techniques still suffer from a dilemma for reasoning designs, ineffective reward, and visual noise. To address these issues, we introduce UI-AGILE, a comprehensive framework enhancing GUI agents at both the training and inference stages. For training, we propose a suite of improvements to the Supervised Fine-Tuning (SFT) process: 1) a Continuous Reward function to incentivize high-precision grounding; 2) a “Simple Thinking” reward to balance planning with speed and grounding accuracy; and 3) a Cropping-based Resampling strategy to mitigate the sparse reward problem and improve learning on complex tasks. For inference, we present Decomposed Grounding with Selection, a novel method that dramatically improves grounding accuracy on high-resolution displays by breaking the image into smaller, manageable parts. Experiments show that UI-AGILE achieves the state-of-the-art performance on two benchmarks ScreenSpot-Pro and ScreenSpot-v2. For instance, using both our proposed training and inference enhancement methods brings 23% grounding accuracy improvement over the best baseline on ScreenSpot-Pro.
@@ 1 Introduction
Figure 1: An overview of UI-AGILE, illustrating its training and inference pipelines. (1) Left: The training pipeline enhances the RFT process with our three core contributions: “Simple Thinking”, continuous grounding reward and cropping-based resampling. Continuous grounding reward being zero would result in crop-based resampling. (2) Right: The inference pipeline contains our proposed decomposed grounding with selection.
@@
Prior methods for GUI agents mostly rely on Supervised Fine-Tuning (SFT), requiring a large amount of human-annotated or synthesized data for teaching the agent how to plan its actions and grounding. (Qin et al. 2025; Cheng et al. 2024; Gou et al. 2025; Wu et al. 2025; Xu et al. 2024; Lin et al. 2025). Recently, Reinforcement Fine-Tuning (RFT) has emerged as an efficient and scalable way to instill reasoning abilities in MLLMs (Chen et al. 2025; Chen, Luo, and Li 2025; Liu et al. 2025b; Peng et al. 2025; Meng et al. 2025). For instance, UI-R1 (Lu et al. 2025) and GUI-R1 (Luo et al. 2025) apply RFT to enhance GUI agents and achieve considerable improvements compared to previous approaches.
@@
To address the above problems, we propose UI-AGILE, a framework aiming at improving both the RFT and the inference stages of GUI agents. The contributions of UI-AGILE can be summarized as follows:
  • To overcome P1, UI-AGILE applies a “Simple Thinking” strategy (Sec. 3.1) that employs reasoning through thoughts with appropriate lengths. UI-AGILE operationalize “Simple Thinking” through a specialized reward function. “Simple Thinking” effectively balances the improvement of both the core grounding task and the prediction of non-grounding action types.
  • To tackle P2, UI-AGILE harnesses a design of continuous grounding reward (Sec. 3.2) for the RFT stage instead of using the common binary reward to incentivize more precise localization to the target’s center. Furthermore, UI-AGILE employs cropping-based resampling (Sec. 3.3) to dynamically adjust the difficulty of training samples to avoid ineffective training with zero reward.
  • To solve P3, UI-AGILE uses a visual noise reduction method termed decomposed grounding with selection (Sec. 3.4). It decomposes a high-resolution screenshot into multiple sub-images, generates candidate elements on each, and finally uses a Vision-Language Model (VLM) to “adjudicate” the best match. This approach significantly improves the agent’s grounding accuracy on high-resolution displays during inference.
@@
Overall, on two benchmarks ScreenSpot-Pro and ScreenSpot-v2, our methods achieve the state-of-the-art performance. For instance, using both our proposed training and inference enhancement methods brings 23% grounding accuracy improvement over the best baseline on ScreenSpot-Pro.
@@ 2 Related Work
Recently, reinforcement learning (RL) techniques for training large models have gained significant momentum. The focus has shifted from traditional policy optimization algorithms like PPO (Schulman et al. 2017), towards alignment-centric methods such as DPO (Rafailov et al. 2023), and more recent rule-based algorithms like GRPO (Shao et al. 2024). These algorithms have achieved remarkable success in enhancing the reasoning capabilities of large models on complex tasks, with models like OpenAI O1 (Jaech et al. 2024) and DeepSeek-R1 (DeepSeek-AI et al. 2025) setting new standards in mathematics and code generation. The efficacy of these approaches prompted their rapid extension into the multimodal domain (Chen et al. 2025; Chen, Luo, and Li 2025; Liu et al. 2025b; Peng et al. 2025; Meng et al. 2025).
@@
The field of GUI agents has evolved rapidly (Wang et al. 2024; Li and Huang 2025). Following early works like CogAgent (Hong et al. 2024) and SeeClick (Cheng et al. 2024), recent studies rely on SFT to operate directly on visual inputs. Show-UI (Lin et al. 2025) innovates on visual processing efficiency. OS-Atlas (Wu et al. 2025), UGround (Gou et al. 2025) and Aria-UI (Yang et al. 2024) propose novel, large-scale pipelines to collect and synthesize millions of GUI agent trajectories, significantly improving model generalization. Aguvis (Xu et al. 2024) introduced a two-stage training process that explicitly uses VLM-generated Chain-of-Thought (CoT) data to teach planning and reasoning. JEDI (Xie et al. 2025) constructs a refusal part by mismatching existing instructions with unrelated screenshots. Standing out in complexity and scale, UI-TARS (Qin et al. 2025) utilizes the largest dataset and the most intricate training pipeline, which involves SFT and DPO on human annotated CoT data to improve performance. This data-intensive scaling has motivated a shift towards more effective RL paradigms, first explored by UI-R1 (Lu et al. 2025) and GUI-R1 (Luo et al. 2025). InfiGUI-R1 (Liu et al. 2025a) employs Spatial Reasoning Distillation and RL to enhance planning and error recovery capabilities. GUI-G1 (Zhou et al. 2025) leverages Hit-based reward and IoU-based reward for improving GUI agents.
@@ 3 Our Framework UI-AGILE
At training time, UI-AGILE adopts a new design for the reward function, which consists of a “Simple Thinking” reward for efficient reasoning (Sec. 3.1) and a continuous grounding reward (Sec. 3.2) for precise localization. UI-AGILE further uses cropping-based resampling (Sec. 3.3), a novel strategy designed to overcome the sparse reward issue. The model generates multiple responses from an image and instruction, which are evaluated by reward functions, including “Simple Thinking” reward and continuous grounding reward. If a training sample proves too difficult (i.e., receive a grounding reward of zero for all generated responses), the image will be cropped to simplify the task, and the model resamples new responses from this modified input.
@@
GRPO algorithm (Shao et al. 2024) is widely recognized as a powerful technique for instilling complex reasoning abilities in LLMs, often through rewarding elaborate chains of thought. Lu et al. (2025) posit that excessive reasoning is not essential for GUI grounding and can even be detrimental. However, the complete role of GUI agents extends beyond mere grounding, since deciding next action (e.g., click or type) inherently demands a foundational level of reasoning.
@@
The total reward for the thought process, $R_{think}$ is formally defined as:
@@
“Simple Thinking” defines an ideal range where the reward is maximized, encouraging thoughts that are neither too brief (“under-thinking”) nor too verbose (“over-thinking”). Outside this ideal range, it uses cosine functions for smooth degradation down to a reward of zero at the absolute bounds. This smooth, non-linear penalty provides a more informative and stable learning signal for RL than a hard cliff, gently guiding the model toward the desired length. Furthermore, the additional bonus for syntactically complete thoughts discourages incomplete reasoning, thereby ensuring better training stability.
@@ 3.2 Continuous Grounding Reward for Precise Localization (P2)
Prior works (Lu et al. 2025; Luo et al. 2025) shift the focus of GUI agent evaluation from traditional object grounding (i.e., IoU) to the precision of the action coordinate. To this end, they typically employ a simple binary reward for localization: a reward of 1 for a correct prediction (e.g., inside the target radius) and 0 otherwise.
However, this binary reward is insufficient for high-precision control, as it treats all successful predictions equally. For instance, a point on the target’s edge receives the same reward as one at its center. This non-discriminatory feedback misguides the model to learn an element’s boundaries rather than its semantic core.
@@ 3.3 Cropping-Based Resampling for Sparse Reward Mitigation (P2)
During GRPO training, GUI agents often face the sparse reward challenge, particularly on complex tasks. When the model consistently fails to place its prediction within the correct bounding box for a given screenshot, it receives a persistent reward of 0. The lack of positive signals can lead to training stagnation, and difficult samples cannot contribute to model improvement.
Inspired by curriculum learning (Bengio et al. 2009), a training technique that trains the model on examples of increasing difficulty to ease the training process, we propose a cropping-based resampling method for mitigating the sparse reward issue. It acts as a dynamic difficulty adjustment mechanism during the training of UI-AGILE. If a task sample yields zero reward over multiple generations, we hypothesize that the task sample is currently too difficult for the model. Then, we reduce the task’s complexity by cropping the original screenshot. The cropping is generated such that it is smaller than the original view but still fully contains the ground-truth bounding box of the target element.
@@
Input: $Image$, $Bbox$, $scalingFactorf$
@@
12:   if $Bbox$ is contained within $Coord_{crop}$ then
@@
19:return $Set$
Modern electronic devices feature high-resolution displays (e.g., 3840x2160), which, when converted into tokens for a VLM, can result in an overwhelmingly large input sequence (e.g., over 10,000 tokens). We hypothesize that a significant portion of these tokens represent irrelevant background information, acting as noise that can distract the GUI agent and degrade its grounding accuracy, i.e., the visual noise issue (P1) mentioned in Sec. 1.
To validate this hypothesis, we conduct a preliminary experiment on ScreenSpot-Pro (Li et al. 2025). We apply the cropping method in Sec. 3.3, but for the purpose of creating a controlled test environment. For each original screenshot, we crop it to 1024x1024, ensuring the ground-truth bounding box is contained within the frame. On this new dataset, the grounding accuracy of UGround-V1-7B (Gou et al. 2025) shows a significant improvement from 31.6 to 56.0 w.r.t grounding accuracy, verifying our hypothesis.
@@ 3.4 Decomposed Grounding with Selection for Visual Noise Reduction (P3)
    Selection: For the final selection stage, we prompt the VLM with the user’s instruction, the candidate element image, and a direct question asking whether the image match the instruction. Then use the model’s output logit for the “Yes” token as a direct relevance score for the candidate. The candidate with the highest “Yes” score is chosen as the final answer, and its corresponding coordinates are remapped to the original screenshot. This QA-based scoring allows the model to perform deeper contextual prediction.
This process directly benefits from our continuous reward function (Sec. 3.2), as it trains the model to predict points closer to the center of target elements, leading to higher-quality extracted images and thus a more accurate final selection.
Analysis of Inference Cost
@@
While the decoding process runs for each sub-image, the small number of output tokens (compared to image tokens) per run ensures that the cumulative decoding cost will not raise much.
@@
Overall, the cost of applying decomposed grounding with selection is low. In addition to the above analysis, we provide results on actual running time in Sec. 4.6. Crucially, we believe this overhead could be eliminated or even reversed with future optimizations in inference engines tailored for this “many small requests” workload, potentially making our high-accuracy method faster than the baseline.
@@ 4.1 Implementation Details
We collect data related to GUI tasks from multiple open-source datasets, including UI-R1 (Lu et al. 2025), GUI-R1 (Luo et al. 2025), Aguvis (Xu et al. 2024) and Grounding-R1 (Yang et al. 2025). We filter them using OmniParser (Wan et al. 2024) following Grounding-R1 (Yang et al. 2025). We randomly sample about 9k examples to train UI-AGILE-3B and UI-AGILE-7B.
Training Details.
We use the trl framework (von Werra et al. 2020) to implement the cropping-based resampling strategy and reward functions. The sampling process is attempted 4 times at most and is bypassed entirely if the bbox’s dimensions exceed the target crop size. Following prior works (Lu et al. 2025; Luo et al. 2025; Liu et al. 2025a; Zhou et al. 2025), we use Qwen2.5-VL-3B [^1] https://huggingface.co/Qwen/Qwen2.5-VL-3B-Instruct and Qwen2.5-VL-7B [^2] https://huggingface.co/Qwen/Qwen2.5-VL-7B-Instruct as base models. Tab. 1 provides the default hyperparameters where cropping factor is the width and height ratio of new attempted image and last attempted image.
Hyperparameter Value
learning rate from 1e-06 to 4.36e-10
num generations 8
num train epochs 2
per device train batch size 4
cropping factor 0.6
gradient accumulation steps 4
Table 1: Default Hyperparameters
Inference Details.
For decomposed grounding with selection, the input image is divided into four sub-images scaling to 60% of the original dimensions, with adjacent sub-images overlapping by 10% of the original image’s width and height. In the element image extraction stage, we define the element’s area by creating a simple bounding box centered on the predicted point with the width and height equal to 14% of the sub-image’s width and height. We have also explored a more sophisticated approach using OmniParser to refine this bounding box. However, it does not improve performance and increases the inference overhead. In the selection stage, we use Qwen2.5VL-7B-instruct11footnotemark: 1 to choose the final answer.
@@ 4.2 Grounding Capability Evaluation
Effectiveness of the proposed inference enhancement.
As shown in Tab. 2, our decomposed grounding with selection method shows significant improvements on the challenging ScreenSpot-Pro benchmark. It provides a universal and substantial performance boost across all tested models, regardless of their original training paradigm (SFT or RFT). For instance, it elevates the average score of OS-Atlas-7B from 18.9 to 33.1 (+75.1%), and boosts Aguvis-7B from 20.4 to 36.5 (+78.9%). The consistent improvement validates the effectiveness of decomposed grounding with selection and its high applicability as a plug-and-play inference enhancement.
Effectiveness of the proposed training enhancement.
Besides, Tab. 2 shows that our core UI-AGILE-3B and UI-AGILE-7B models, even without decomposed grounding, establish a new state-of-the-art among 3B and 7B models on ScreenSpot-Pro. Trained on only 9K examples with 2 epochs, they (37.9 for 3B and 44.0 for 7B) surpass other RFT-based models like UI-R1-E (33.5), InfiGUI-R1-3B and GUI-R1-7B (32.1). UI-AGILE-7B even outperforms the much larger model UI-TARS-72B (38.1) trained on approximately 50 billion tokens. On the ScreenSpot-v2 benchmark (Tab. 3), our UI-AGILE-7B also achieves state-of-the-art grounding accuracy with an average score of 92.1. The above results demonstrate the effectiveness of our proposed “Simple Thinking” reward, continuous grounding reward, and cropping-based resampling for improving the training of GUI agents.
Overall, as shown in Tab. 2, using both our proposed training and inference enhancement methods (UI-AGILE-7B + Decomposed Grounding) brings 23% grounding accuracy improvement over the best baseline (JEDI-7B) on ScreenSpot-Pro.

### v3

@@ Abstract
The emergence of Multimodal Large Language Models (MLLMs) has driven significant advances in Graphical User Interface (GUI) agent capabilities. Nevertheless, existing GUI agent training and inference techniques still suffer from a dilemma for reasoning designs, ineffective reward, and visual noise. To address these issues, we introduce UI-AGILE for enhancing GUI agents at both training and inference. For training, we propose a suite of improvements to the Supervised Fine-Tuning (SFT) process: 1) a continuous reward function to incentivize high-precision grounding; 2) a “Simple Thinking” reward to balance planning with speed and grounding accuracy; and 3) a cropping-based resampling strategy to mitigate the sparse reward problem and improve learning on complex tasks. For inference, we present decomposed grounding with selection to dramatically improve grounding accuracy on high-resolution displays by breaking the image into smaller, manageable parts. Experiments show that UI-AGILE achieves the state-of-the-art grounding performance on two benchmarks ScreenSpot-Pro and ScreenSpot-v2 while it also exhibits strong general agent capabilities. For instance, using both our training and inference enhancement methods brings 23% grounding accuracy improvement over the best baseline on ScreenSpot-Pro. We provide the code in https://github.com/KDEGroup/UI-AGILE.
@@ 1 Introduction
Figure 1: An overview of UI-AGILE. (1) Left: The training stage is enhanced with our three core contributions: “Simple Thinking”, continuous grounding reward and cropping-based resampling. Continuous grounding reward being zero would result in crop-based resampling. (2) Right: The inference stage uses our proposed decomposed grounding with selection.
@@
Prior methods for GUI agents mostly rely on Supervised Fine-Tuning (SFT), requiring a large amount of human-annotated or synthesized data for teaching agents how to plan its actions and grounding (Gou et al. 2025; Wu et al. 2025; Lin et al. 2025). Recently, Reinforcement Fine-Tuning (RFT) has been proposed to enhance GUI agents (Lu et al. 2025; Luo et al. 2025), and achieves considerable improvements compared to previous approaches.
@@
To address the above problems, we propose UI-AGILE, a framework aiming at improving both the RFT and the inference stages of GUI agents. Its contributions are as follows:
  • To overcome P1, UI-AGILE applies a “Simple Thinking” strategy that employs reasoning through thoughts with appropriate lengths. “Simple Thinking” effectively balances the improvement of both the core grounding task and the prediction of non-grounding action types.
  • To tackle P2, UI-AGILE harnesses a design of continuous grounding reward for the RFT stage to incentivize more precise localization to the target’s center. Furthermore, UI-AGILE employs cropping-based resampling to dynamically adjust the difficulty of training samples to avoid ineffective training with zero reward.
  • To solve P3, UI-AGILE uses a visual noise reduction method termed decomposed grounding with selection. It decomposes a high-resolution screenshot into multiple sub-images, generates candidate elements on each, and finally uses a Vision-Language Model (VLM) to “adjudicate” the best match. This approach significantly improves the agent’s grounding accuracy on high-resolution displays during inference.
@@ 2 Related Work
Recently, reinforcement learning (RL) for training large models has gained significant momentum. The focus has shifted from traditional policy optimization algorithms like PPO (Schulman et al. 2017), towards alignment-centric methods such as DPO (Rafailov et al. 2023), and more recent rule-based algorithms like GRPO (Shao et al. 2024). These algorithms have achieved remarkable success in enhancing the reasoning capabilities of large models on complex tasks, with models like OpenAI O1 (Jaech et al. 2024) and DeepSeek-R1 (DeepSeek-AI et al. 2025). The efficacy of these approaches promotes their rapid extension into the multimodal domain (Chen et al. 2025; Chen, Luo, and Li 2025; Liu et al. 2025b; Peng et al. 2025; Meng et al. 2025).
@@
The field of GUI agents has evolved rapidly (Wang et al. 2024; Li and Huang 2025). Following early works like CogAgent (Hong et al. 2024) and SeeClick (Cheng et al. 2024), most studies rely on SFT to operate directly on visual inputs. Show-UI (Lin et al. 2025) innovates on visual processing efficiency. OS-Atlas (Wu et al. 2025), UGround (Gou et al. 2025) and Aria-UI (Yang et al. 2024) propose novel, large-scale pipelines to collect and synthesize millions of GUI agent trajectories, significantly improving model generalization. Aguvis (Xu et al. 2024) introduced a two-stage training process that explicitly uses VLM-generated Chain-of-Thought (CoT) data to teach planning and reasoning. JEDI (Xie et al. 2025) constructs a refusal part by mismatching existing instructions with unrelated screenshots. Standing out in complexity and scale, UI-TARS (Qin et al. 2025) utilizes the largest dataset and the most intricate training pipeline, which involves SFT and DPO on human annotated CoT data to improve performance. This data-intensive scaling has motivated a shift towards more effective RL paradigms, first explored by UI-R1 (Lu et al. 2025) and GUI-R1 (Luo et al. 2025). InfiGUI-R1 (Liu et al. 2025a) employs Spatial Reasoning Distillation and RL to enhance planning and error recovery capabilities. GUI-G1 (Zhou et al. 2025) leverages Hit-based reward and IoU-based reward for improving GUI agents.
@@ 3 Our Framework UI-AGILE
At training time, the model generates multiple responses from an image and instruction, which are evaluated by reward functions, including “Simple Thinking” reward for efficient reasoning (Sec. 3.1) and continuous grounding reward (Sec. 3.2) for precise localization. If a training sample proves too difficult (i.e., receive a grounding reward of zero for all generated responses), the image will be cropped to simplify the task, and the model resamples new responses from this modified input.
@@
Lu et al. (2025) posit that excessive reasoning is not essential for GUI grounding and can even be detrimental. However, the complete role of GUI agents extends beyond mere grounding, and deciding next action (e.g., click or type) inherently demands a foundational level of reasoning.
@@
The reward $R_{think}$ for the thought process is defined as:
@@
“Simple Thinking” defines an ideal range where the reward is maximized, encouraging thoughts that are neither too brief (“under-thinking”) nor too verbose (“over-thinking”). Outside this ideal range, it uses the cosine function for smooth degradation down to a reward of zero at the absolute bounds. This smooth, non-linear penalty provides a more stable learning signal for RL than a hard cliff. Furthermore, the additional bonus for syntactically complete thoughts discourages incomplete reasoning, thereby ensuring better training stability.
@@ 3.2 Continuous Grounding Reward for Precise Localization (P2)
Unlike general visual grounding tasks which compute the IoU between the predicted bounding box and the ground truth box, GUI agents usually predict a click point and receive a simple binary reward in rule-based RFT (Lu et al. 2025; Luo et al. 2025): a reward of 1 for a correct prediction (e.g., inside the ground truth box) and 0 otherwise.
However, this binary reward is insufficient for high-precision control, as it equally rewards clicks on an element’s edge and at its center. This non-discriminatory feedback misguides the model to learn an element’s boundaries rather than its semantic core.
@@ 3.3 Cropping-Based Resampling for Sparse Reward Mitigation (P2)
During GRPO training, GUI agents often face the sparse reward challenge, particularly on complex tasks. When the model fails to place its prediction within the correct bounding box for a given screenshot over all generations, it receives reward of 0. The lack of positive signals can lead to training stagnation, and difficult samples cannot contribute to model improvement.
Inspired by curriculum learning (Bengio et al. 2009), which presents training examples in an easy-to-hard progression, we propose Cropping-based Resampling. This method acts as a dynamic difficulty adjustment mechanism: if a sample yields zero reward over all generations, we deem it too difficult and reduce its complexity by cropping the screenshot. The resulting smaller crop is guaranteed to fully contain the ground-truth bounding box.
@@ 3.4 Decomposed Grounding with Selection for Visual Noise Reduction (P3)
Modern electronic devices feature high-resolution displays (e.g., 3840x2160), which, when converted into tokens for a VLM, can result in an overwhelmingly long sequence (e.g., over 10,000 tokens). We hypothesize that a significant portion of these tokens represent irrelevant background information, i.e., the visual noise issue P1 mentioned in Sec. 1.
To validate this hypothesis, we conduct a preliminary experiment on ScreenSpot-Pro (Li et al. 2025). We apply the cropping method in Sec. 3.3, but for the purpose of creating a controlled test environment. For each original screenshot, we crop it to 1024x1024, ensuring the ground-truth bounding box is contained within the frame. On this new dataset, the grounding accuracy of UGround-V1-7B (Gou et al. 2025) shows a significant improvement from 31.6 to 56.0, verifying our hypothesis.
@@
Input: $Image$, BboxBbox, scalingFactorfscalingFactorf
@@
12:   if BboxBbox is contained within $Coord_{crop}$ then
@@
19:return SetSet
@@
    Selection: We use a VLM to score each candidate image based on a direct “Yes/No” question asking if it matches the user’s instruction (see Appendix A for the prompt). The output logit for “Yes” serves as the relevance score. The coordinates of the highest-scoring candidate are then remapped to the original screenshot. This QA-based process enables deeper contextual prediction.
This process directly benefits from the continuous reward function, as it trains the model to predict points closer to the center of target elements, leading to higher-quality extracted images and thus a more accurate final selection.
Analysis of Inference Cost.
@@
While the decoding process runs for each sub-image, the small number of output tokens (compared to image tokens) per run ensures that the cumulative decoding cost will not increase much.
@@
Overall, the cost of applying decomposed grounding with selection is low. In addition to the above analysis, we provide results on actual running time in Sec. 4.5. Crucially, we believe this overhead could be eliminated or even reversed with future optimizations in inference engines tailored for this “many small requests” workload.
@@ 4.1 Implementation Details
We collect data from multiple open-source datasets, including UI-R1 (Lu et al. 2025), GUI-R1 (Luo et al. 2025), Aguvis (Xu et al. 2024) and Grounding-R1 (Yang et al. 2025). We filter them using OmniParser (Wan et al. 2024) following Grounding-R1 (Yang et al. 2025). We randomly sample about 9k examples to train UI-AGILE.
Details of Environment, Implementation and Default Hyperparameters.
Due to space limit, we include the details of environment, implementation and default hyperparameters in Appendix A.
@@ 4.2 Grounding Capability Evaluation
Effectiveness of our inference enhancement.
As shown in Tab. 1, decomposed grounding with selection shows significant improvements on ScreenSpot-Pro. It provides a universal and substantial performance boost across all tested models, regardless of their original training paradigm (SFT or RFT). For instance, it elevates the average score of OS-Atlas-7B from 18.9 to 33.1 (+75.1%), and boosts Aguvis-7B from 20.4 to 36.5 (+78.9%). The consistent improvement proves the effectiveness of decomposed grounding with selection and its high applicability as a plug-and-play inference enhancement.
Effectiveness of our training enhancement.
Besides, Tab. 1 shows that UI-AGILE-3B and UI-AGILE-7B models, even without decomposed grounding, establish a new state-of-the-art among 3B and 7B models on ScreenSpot-Pro. Trained on only 9K examples with 2 epochs, they (37.9 for 3B and 44.0 for 7B) surpass other RFT-based models like UI-R1-E (33.5), InfiGUI-R1-3B and GUI-R1-7B (32.1). UI-AGILE-7B even outperforms the much larger model UI-TARS-72B (38.1) trained on approximately 50 billion tokens. On the ScreenSpot-v2 benchmark (Tab. 2), our UI-AGILE-7B also achieves state-of-the-art grounding accuracy with an average score of 92.1. The above results demonstrate the effectiveness of our proposed “Simple Thinking” reward, continuous grounding reward, and cropping-based resampling for improving the training of GUI agents.
Overall, as shown in Tab. 1, using both our training and inference enhancements (UI-AGILE-7B + Decomposed Grounding) brings 23% grounding accuracy improvement over the best baseline (JEDI-7B) on ScreenSpot-Pro.